8. The activity of a radioactive sample fell from 2400Bq to 75 Bq over period of 24 hours. How long is the isotope’s half life?


1
Expert's answer
2021-03-26T11:26:36-0400

We know that change in activity depends on the half-life and time:


"A=\frac{A_0}{2^{t/T}},\\\space\\
\frac{A_0}{A}=2^{t/T},\\\space\\
T=\frac{t\ln2}{\ln(A_0/A)}=4.8\text{ h}."